OpenStudy (anonymous):

A manufacturer is using a cubic box with edges of 8.23 cm each for his product. Estimate the volume of the box using whole cm.

OpenStudy (anonymous):

Since you are asked to estimate the volume, not solve it exactly, you should make an estimate of the side length that will make the math easy...

OpenStudy (anonymous):

Instead of using sides 8.23 cm, maybe you could estimate that as just 8 cm.

OpenStudy (anonymous):

V=x^3 x=8.23 estimate x=8 so 8^3 8*8*8=